Identify the means and extremes, then, find the missing terms.
4:5 = 24:x

Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39626)
Answer by greenestamps(13203)



I am not a fan of this terminology; but....

In the proportion A:B = C:D, the extremes are the numbers "to the extreme outsides of the proportion" -- A and D. The means are then the other two numbers -- closest to the equal sign, B and C.

The unpleasant rule is then "the product of the means is equal to the product of the extremes": BC = AD.

ANSWERS:

(1) In the given proportion, the means are 5 and 24; the extremes are 4 and x.

(2) 5*24 = 120 = 4*x --> x=30